Show that every open interval is an open set


1
Expert's answer
2021-02-04T07:40:49-0500

Let "(i,j)" be an open interval. Let "a \in (i, j)" . We show that, we can find "\epsilon > 0 \ni" "N(a, \epsilon) \sub (i,j)"

Since "a \in (i,j)" we have that "i<a<j" this implies that "a-i > 0" and "j-a > 0"

Set "\epsilon = min (a-i, j-a)"

Suppose further that "b \in N(a, \epsilon)" this implies that "|b-a| < \epsilon"

"- \epsilon < b - a< \epsilon \\
a - \epsilon < b < a + \epsilon"

And since "\epsilon" is as defined we have that "a - \epsilon \geq i \\
a + \epsilon \leq j"

This implies that "b \in (i, j)"

Hence "N(a, \epsilon ) \sub (i,j)"